# On the reconstruction limits of complex networks

### v1

@@ IV.3 Role of evidence-based model selection
where $P(X_{i}=x)$ is the evidence of model $M_{i}$ for $x$. If $B_{M_{1},M_{2}}(x)>1$, $M_{1}$ is better supported by the data $x$ than $M_{2}$. From an information-theoretic perspective, the minimization of the evidence CE—which is equivalent to maximizing the evidence—can be shown to be a necessary condition for finding the TDG model. Indeed, for a reconstruction model $M$ with evidence $P(\hat{X}=x)$, the evidence CE is expressed as
@@
The problem of evidence-based model selection is the computation of the evidence itself which is often intractable in practice. In fact, this is the case for most graph models, where the evaluation of the evidence requires graph enumeration. This problem also arises in the evaluation of both the information gain and the mutual information. Fortunately, the same numerical techniques can be used to evaluate the evidence and the information gain simultaneously, as the two are related to each other. In Ref. Murphy et al. (2024), we showed that variational mean-field methods provide efficient approximations for both the mutual information and the evidence. The same techniques are used here (see Sec. G).
Model selection is crucial for the validity of the reconstruction index as a proxy of reconstruction performance. Suppose we observed a set of observations $\left\{x_{1},x_{2},...\right\}$ and used the reconstruction model $M^{*}$. The empirical average of the information gain of some model $M$ converges to
$\mathbb{E}_{X^{*}}\quantity[\mathcal{I}_{M}(X^{*})]$ (31)
@@
Now, assume for a moment that $M$ in fact maximizes the expected log-evidence. This implies, as we mentionned before, that $X^{*}$ and $X$ are equal in distribution. In turn, Eq. (31) becomes equal to the mutual information for model $M$, i.e., $I(X;G)$, which we recall is the reconstruction limit of $M$. This means that Eq. (31) converges to the reconstructability and that ultimately the reconstruction index naturally extends the concept of reconstruction limit to real systems observed only through data. When the reconstruction model does not minimize the evidence CE, the picture becomes more nuanced, as shown in Fig. 5. As the evidence CE decreases, the correlation between the reconstruction index and the posterior loss increases. Our ability to identify the reconstruction limit without knowing the true model or network structure is therefore as good as the reconstruction model’s ability to describe the data. This key conceptual observation leads us to conclude that we can indeed leverage the reconstruction index as a proxy for assessing the reconstructability of real networks, provided it is interpreted in conjunction with the posterior loss, as in Fig. 5(a).
V Network reconstructability in empirical networks
@@
Fig. 6: Reconstruction from spontaneous neuronal activity in the mouse brain Steinmetz et al. (2019); Stringer et al. (2019): (a) Raster plot of the 1462 monitored neurons, (b) reconstruction of the probe network using different reconstruction models and (c) reconstructability diagram. In panel (a), the neurons are ordered by the probe they were measured from. Each spike is represented in blue. Panel (b) shows the posterior average network projected onto the probes, as predicted by each reconstruction model where rows correspond to different graph models (see Appendix A), and columns to different dynamics models (see Table 2). The color of an edge connecting two probes shows the absolute number of edges and the thickness indicates the average proportion among all the edges. The size of the probe nodes is proportional to the number of neurons monitored by the probe, and the color indicates the measured number of spikes. The node locations correspond to the actual probe locations in the mouse brain obtained from Steinmetz et al. (2019). The reconstruction index as a function of the model log evidence is shown in panel (c), comparing the different models. Small markers are estimated by a single Markov chain and large markers are the average of these estimations. In these experiments, the parameters of the graph prior and and likelihood are inferred jointly with the graph. For additional details about the inference procedure, we refer to Appendix J.
@@
V.1 Reconstruction from empirical neuronal spiking data
@@
We consider the spontaneous activity of 1462 neurons from a mouse brain recorded over 20 minutes using eight neopixel probes [Fig. 6(a)] Stringer et al. (2019). Starting from the recorded spike times of the neurons, indicating when they fire, we create a binary time series of the activity of each neuron. In these binary time series, a ‘1’ marks the moments when a neuron is fired, and a ‘0’ when it is not. For more detail regarding our data processing procedure, see Appendix J.1. Then, we infer the network using a variety of Bayesian reconstruction models. As a result, we get the reconstructed networks shown in Fig. 6(b), where we aggregated the edges connecting the neurons of all pairs of probes, thus illustrating how they interact. Finally, we compute the log evidence for each model and select the model with the highest one (see Appendix G for details on the evidence estimation). In doing so, we extract the reconstruction index that is, in principle, closest to the reconstruction limit, given the set of considered models. Figure 6(c) shows a diagram of both measures for each model.
@@
In our example, the analyzed reconstruction models are combinations of time series likelihoods—SIS, Cowan and Glauber—and graph priors—the ER model, the configuration model with a uniform degree sequence hyperprior (UCM) and the stochastic block model (SBM). The ER model, having a uniform distribution, is the most entropic model, followed by the SBM and the CM. Additional details about the graph prior and about these reconstruction models are given in Appendices A and J, respectively. Of course, these models oversimplify the observed neuronal activity. Moreover, certain critical factors are not captured in the current dataset, such as the latency and deactivation rates of neuronal activity, as well as the substantial number of neurons undetected by the probes, which could contribute as input currents to the modeled neurons. In addition, the lack of detailed connectomic data for such small brain regions prevents a rigorous assessment of the accuracy reconstructed probe network. The purpose of this analysis is therefore not to perform the most accurate reconstruction but to illustrate the complete procedure as well as the results it generates.
@@
Among the considered models, the SIS model with a SBM prior is the one achieving the highest log evidence. As detailed in Appendix J, the inferred SIS model reproduces many of the statistics featured in the data, suggesting that its corresponding reconstruction index of approximately $67\%$ is reasonable. Considering the tendency observed Fig. 6(c), one could conjecture that more detailed neuronal models, better suited for these data and with potentially higher log evidence, would yield reconstruction indices even lower than $67\%$. In the following section, we circumvent the limitations of the neuronal activity data by transitioning to a controlled setting where synthetic activity data is used to reconstruct empirical networks.
V.2 Reconstruction of empirical graphs from synthetic activity data
@@
Table 1: Negative log probability—i.e., $-\log P(G=g^{*})$—of the graphs considered in Fig. (7) using the Erdős-Rényi (ER) model, the configuration model with uniform degree sequence prior (UCM), the configuration model with given degree sequence (CM) and the stochastic block model (SBM).
@@
The reconstruction limit, as measured by $\Psi_{M}$, changes as a function of the graph prior: If information about $g^{*}$ is a priori given, this same information cannot be reconstructed from realizations of $X^{*}$. Thus, it is no longer taken into account in $\Psi_{M}$ which by construction factors out the contributions of the prior. This intuition can be mathematically studied if we let the prior put more and more weight on $g^{*}$. In Appendix I, we prove that as the graph generative model converges to a Kronecker delta distribution, i.e.,
@@
VI Conclusion
@@
Our information-theoretical framework reveals an inherent network reconstruction limit, closely tied to the true generating process of the observed data. The reconstruction limit is analogous to the detectability limit in community detection Decelle et al. (2011); Ghasemian et al. (2016); Young et al. (2017) in that it is algorithm independent. We find that this limit is expressed in terms of the reconstructability—a normalized version of the mutual information between the graph and the data of the true data generating process. While a small reconstructability implies a bad performance regardless of the reconstruction model, a high reconstructability implies that good performance can be achieved with the appropriate model.
@@
Our framework is general and can be extended to real modeling settings, where the data is limited and the reconstruction model is unknown. Using the same principles, we defined the reconstruction index, analogous to the reconstructability, that is also data-dependent and can be used as a proxy of the reconstruction performance. When coupled with evidence-based model selection, the reconstruction index is an appropriate performance measure, even when the graph is unknown. This further emphasizes the importance of model quality in network reconstruction Peel et al. (2022).
Finally, we illustrated some applications of our framework using real networks and real time series data. We showed how our framework works in the reconstruction of spiking neural networks. Although the reconstructability is based on random variables and ensembles, we showed that our framework can be reliably used on single realizations of graphs. Our work offers an additional step towards more principled approaches to network reconstruction.
Acknowledgments
Appendix A Graph priors
@@
In the paper, we use different random graph models as graph priors for Bayesian network reconstruction. These models are undirected and unweighted, and may include self-loops and multi-edges, although our general framework is not restricted to these assumptions. Indeed, one could consider directed or weighted graphs as well; as long as the set of possible graphs remains countable. We use the adjacency matrix, denoted $\bm{a}$, in order to define the probability distribution of some of these models, where $a_{ij}$ counts the number of edges connecting nodes $i$ and $j$. To simplify the notation, we will sometimes express a graph $g$ directly with its adjacency matrix $g=\bm{a}$ We use the convention that $a_{ii}$ is always a multiple of 2. Below, we describe these priors in more detail.
A.1 Erdős-Rényi model
@@
such that ((${}{n}{k}\right)\kern-3.00003pt\right)=\binom{n+k-1}{k}$ counts the number of possible multisets of size $k$ composed of $n$ different objects—i.e., multi-edges.
Note that the number of edges $E$ must be provided to the ER model, and in the other graph models described below. This means that $\theta=E$ is the hyperparameter of ER graph prior and that $E$ should be inferred. We use the prior $P(E)$ to weigh in the number of edges. In most of our experiments, the number of edges is fixed to a specific value $e^{*}$, meaning that $P(E=e)=\delta_{e,e^{*}}$, where $\delta_{m,n}$ is the Kronecker delta function. However, in Sec. V.1, as $E$ is unknown in this case, we use a geometric prior of the form
@@
where $\bar{\lambda}$ is a parameter that fixes the expected number of edges. See also Appendix J for further detail about the complete inference procedure.
A.2 Configuration model
@@
$P(\bm{k}=\bm{\kappa})=\delta_{\bm{\kappa},\bm{\kappa}^{*}}\,.$ (38)
@@
A.3 Stochastic block model
The stochastic block model (SBM), in its microcanonical version Peixoto (2017), closely resembles the ER model, where edges are picked uniformly at random. However, unlike the ER model, each node $i$ is associated with a random block $b_{i}\in\{1,2,\dots,B\}$ and instead the number of edges $e_{rs}=\sum_{ij}a_{ij}\delta_{b_{i},r}\delta_{b_{j},s}$ connecting two blocks $r$ and $s$ is fixed such that there are $E$ edges in total. We summarize the node partition as a tuple $\bm{b}=(b_{i})_{i=1..N}$ and the number of edges between blocks by the edge matrix $\bm{e}=(e_{rs})_{r,s=1..B}$. The blocks are required to be non-empty. Since $B$ is the number of non-empty blocks in $\bm{b}$ and $\bm{e}$ is completely determined by the graph and $\bm{b}$, $\theta=(\bm{b},E)$ are the hyperparameters of the SBM. Consequently, $\bm{b}$ and $E$ must be inferred jointly with the graph. In theory, one could factor the joint prior probability $P(G,E,\bm{b})$ as $P(G|E,\bm{b})P(E)P(\bm{b})$ assuming $E$ and $\bm{b}$ are independent. However, it is more convenient to factor the prior also using $\bm{e}$ and $B$ as intermediate random variables, following

### v2

@@ II.3.3 Role of evidence-based model selection
where $P(X_{i}=x)$ is the evidence of model $M_{i}$ for $x$. If $B_{M_{1},M_{2}}(x)>1$, $M_{1}$ is better supported by the data $x$ than $M_{2}$. From an information-theoretic perspective, the minimization of the evidence CE—which is equivalent to maximizing the evidence—can be shown to be a necessary condition for finding the TDG model. Indeed, for a reconstruction model $M$ with evidence distribution $P(X)$, the evidence CE is expressed as
@@
The problem of evidence-based model selection is the computation of the evidence itself which is often intractable in practice. In fact, this is the case for most graph models, where the evaluation of the evidence requires graph enumeration. This problem also arises in the evaluation of both the information gain and the mutual information. Fortunately, the same numerical techniques can be used to evaluate the evidence and the information gain simultaneously, as the two are related to each other. In Ref. Murphy et al. (2024), we showed that variational mean-field methods provide efficient approximations for both the mutual information and the evidence. The same techniques are used here (see Sec. IV.6).
@@
Model selection is crucial for the validity of the reconstruction index as a proxy of reconstruction performance. Suppose we observed many observations $(x_{1},x_{2},...)$ and used the reconstruction model $M^{*}$. The empirical average of the information gain of some model $M$ converges to $\mathbb{E}_{X^{*}}\quantity[\mathcal{I}_{M}(X^{*})]$. Now, assume for a moment that $M$ in fact maximizes the expected log-evidence. This implies, as we mentioned before, that $X^{*}$ and $X$ are equal in distribution. In turn, the empirical information gain becomes equal to the mutual information for model $M$, i.e., $I(X;G)$, which we recall is the reconstruction limit of $M$. This means that empirical information gain converges to the reconstructability and that ultimately the reconstruction index naturally extends the concept of reconstruction limit to real systems observed only through data.
When the reconstruction model does not minimize the evidence CE, the picture becomes more nuanced, as shown in Fig. 5. As the evidence CE decreases, the correlation between the reconstruction index and the posterior loss increases. Our ability to identify the reconstruction limit without knowing the true model or network structure is therefore as good as the reconstruction model’s ability to describe the data. This key conceptual observation leads us to conclude that we can indeed leverage the reconstruction index as a proxy for assessing the reconstructability of real networks, provided it is interpreted in conjunction with the posterior loss, as in Fig. 5(a).
II.4 Network reconstructability in empirical networks
@@
Fig. 6: Reconstruction from spontaneous neuronal activity in the mouse brain Steinmetz et al. (2019); Stringer et al. (2019): (a) Raster plot of the 1462 monitored neurons, (b) reconstruction of the probe network using different reconstruction models and (c) reconstructability diagram. In panel (a), the neurons are ordered by the probe they were measured from. Each spike is represented in blue. Panel (b) shows the posterior average network projected onto the probes, as predicted by each reconstruction model where rows correspond to different graph models (see Appendix IV.1), and columns to different dynamics models (see Table 2). The color of an edge connecting two probes shows the absolute number of edges and the thickness indicates the average proportion among all the edges. The size of the probe nodes is proportional to the number of neurons monitored by the probe, and the color indicates the measured number of spikes. The node locations correspond to the actual probe locations in the mouse brain obtained from Steinmetz et al. (2019). The reconstruction index as a function of the model log evidence is shown in panel (c), comparing the different models. Small markers are estimated by a single Markov chain and large markers are the average of these estimations. In these experiments, the parameters of the graph prior and and likelihood are inferred jointly with the graph. For additional details about the inference procedure, we refer to Appendix IV.9.
@@
II.4.1 Reconstruction from empirical neuronal spiking data
@@
We consider the spontaneous activity of 1462 neurons from a mouse brain recorded over 20 minutes using eight neopixel probes [Fig. 6(a)] Stringer et al. (2019). Starting from the recorded spike times of the neurons, indicating when they fire, we create a binary time series of the activity of each neuron. In these binary time series, a ‘1’ marks the moments when a neuron is fired, and a ‘0’ when it is not. For more detail regarding our data processing procedure, see Appendix IV.9. Then, we infer the network using a variety of Bayesian reconstruction models. As a result, we get the reconstructed networks shown in Fig. 6(b), where we aggregated the edges connecting the neurons of all pairs of probes, thus illustrating how they interact. Finally, we compute the log evidence for each model and select the model with the highest one (see Appendix IV.6 for details on the evidence estimation). In doing so, we extract the reconstruction index that is, in principle, closest to the reconstruction limit, given the set of considered models. Figure 6(c) shows a diagram of both measures for each model.
@@
In our example, the analyzed reconstruction models are combinations of time series likelihoods—SIS, Cowan and Glauber—and graph priors—the ER model, the configuration model with a uniform degree sequence hyperprior (UCM) and the stochastic block model (SBM). The ER model, having a uniform distribution, is the most entropic model, followed by the SBM and the CM. Additional details about the graph prior and about these reconstruction models are given in Appendices IV.1 and IV.2, respectively. Of course, these models oversimplify the observed neuronal activity. Moreover, certain critical factors are not captured in the current dataset, such as the latency and deactivation rates of neuronal activity, as well as the substantial number of neurons undetected by the probes, which could contribute as input currents to the modeled neurons. In addition, the lack of detailed connectomic data for such small brain regions prevents a rigorous assessment of the accuracy reconstructed probe network. The purpose of this analysis is therefore not to perform the most accurate reconstruction but to illustrate the complete procedure as well as the results it generates.
@@
Among the considered models, the SIS model with a SBM prior is the one achieving the highest log evidence. As detailed in Appendix IV.9, the inferred SIS model reproduces many of the statistics featured in the data, suggesting that its corresponding reconstruction index of approximately $67\%$ is reasonable. Additionally, given that the network contains $1462$ neurons and that the estimated number of edges for this model is approximately $1722$ (see Table 3), the inferred network is sparser than expected (e.g., see Table 2 in Lin et al. (2024)) with an average degree of $2.35$. This suggests that, although our estimation of reconstructability is not close to zero, the inferred network does not seem to account for most of the neural activity. Of course, considering the decreasing tendency observed in Fig. 6(c), more detailed neuronal models—better suited for these data and with potentially higher log evidence—could yield reconstruction indices even lower than $67\%$, with possibly denser inferred networks. In the following section, we circumvent the limitations of the neuronal activity data by transitioning to a controlled setting, where synthetic activity data is used to reconstruct empirical networks.
II.4.2 Reconstruction of empirical graphs from synthetic activity data
@@
Table 1: Negative log probability—i.e., $-\log P(G=g^{*})$—of the graphs considered in Fig. 7 using the Erdős-Rényi (ER) model, the configuration model with uniform degree sequence prior (UCM), the configuration model with given degree sequence (CM) and the stochastic block model (SBM).
@@
The reconstruction limit, as measured by $\Psi_{M}$, changes as a function of the graph prior: If information about $g^{*}$ is a priori given, this same information cannot be reconstructed from realizations of $X^{*}$. Thus, it is no longer taken into account in $\Psi_{M}$ which by construction factors out the contributions of the prior. This intuition can be mathematically studied if we let the prior put more and more weight on $g^{*}$. In Appendix IV.8, we prove that as the graph generative model converges to a Kronecker delta distribution, i.e.,
@@
III Discussion
To what extent is a complex network reconstructible ? Our ability to reconstruct is strongly constrained by the information content of the underlying structure within the data, making perfect reconstruction generally infeasible. The best reconstruction therefore amounts to finding a network that reaches a reconstruction limit, extracting all available information in such a way that no further improvement can be achieved on average.
@@ III Discussion
Our information-theoretic framework characterizes this network reconstruction limit, which is closely tied to the true generating process of the observed data. The reconstruction limit is analogous to the detectability limit in community detection Decelle et al. (2011); Ghasemian et al. (2016); Young et al. (2017) in that it is algorithm independent. We find that this limit is expressed in terms of the reconstructability—a normalized version of the mutual information between the graph and the data of the true data generating process. While a small reconstructability implies a bad performance regardless of the reconstruction model, a high reconstructability implies that good performance can be achieved with the appropriate model.
@@
Our approach is general and can be extended to real modeling settings, where the data is limited and the reconstruction model is unknown. Using the same principles, we defined the reconstruction index, analogous to the reconstructability, that is also data-dependent and can be used as a proxy of the reconstruction performance. When coupled with evidence-based model selection, the reconstruction index is an appropriate performance measure, even when the graph is unknown. This further emphasizes the importance of model quality in network reconstruction Peel et al. (2022).
Finally, we presented different applications of our framework using real networks and real time series data. We showed how to use the reconstruction index on spiking neural networks. Our analysis suggests that the reconstructability of the network formed by the recorded neurons in the mouse brain is approximately 67%, which is consistent with experimental studies of brain networks reporting structure-function couplings of about 40% in humans Baum et al. (2020) and up to 68% in smaller animal models, such as zebrafish Legare et al. (2024). We believe that more work is needed in the inference process, primarily because of the simplicity of the reconstruction models used in this analysis and the incompleteness of the data. A more thorough analysis of such a case study could reveal that certain neuronal activity datasets are insufficient to build a comprehensive picture of the functional activity of the brain. Additionally, although the reconstructability is based on random variables and ensembles, we demonstrate that our framework can be reliably used on single instances of graphs. In this context, the reconstruction index can be used to determine the reconstruction limit, even if the true graph is unknown, as it is shown to correlate strongly with the error between the inferred graphs and the true one.
We envision a future where network reconstruction applications incorporate a reconstructability analysis in their pipeline, such as that presented in this paper. By doing so, the reconstruction index would indicate how informative the reconstructed networks are and perhaps inform us on how they should be used within the said applications. Of course, there is still plenty of work to be done on this front, such as improving the computational methods required to compute the reconstruction index as they do not scale well to large networks, and improving the reconstruction models themselves as we have alluded to earlier. Some of these models might also require modifying our framework, for example in the case of weighted and directed networks. These specific models could prove considerably valuable for the neuroscience community and, more broadly, for complex systems research.
IV Materials and methods
IV.1 Graph priors
@@
In the paper, we use different random graph models as graph priors for Bayesian network reconstruction. These models are undirected and unweighted and may include self-loops and multiedges, although our general framework is not restricted to these assumptions. Indeed, one could consider directed or weighted graphs as well; as long as the set of possible graphs remains countable. We use the adjacency matrix, denoted $\bm{a}$, in order to define the probability distribution of some of these models, where $a_{ij}$ counts the number of edges connecting nodes $i$ and $j$. To simplify the notation, we will sometimes express a graph $g$ directly with its adjacency matrix $g=\bm{a}$ We use the convention that $a_{ii}$ is always a multiple of 2. Below, we describe these priors in more detail.
IV.1.1 Erdős-Rényi model
@@
such that ((${}{n}{k}\right)\kern-3.00003pt\right)=\binom{n+k-1}{k}$ counts the number of possible multisets of size $k$ composed of $n$ different objects—i.e., multiedges.
Note that the number of edges $E$ must be provided to the ER model, and in the other graph models described below. This means that $\theta=E$ is the hyperparameter of ER graph prior and that $E$ should be inferred. We use the prior $P(E)$ to weigh in the number of edges. In most of our experiments, the number of edges is fixed to a specific value $e^{*}$, meaning that $P(E=e)=\delta(e,e^{*})$, where $\delta(m,n)$ is the Kronecker delta function. However, in Sec. II.4.1, as $E$ is unknown in this case, we use a geometric prior of the form
@@
where $\bar{\lambda}$ is a parameter that fixes the expected number of edges. See also Appendix IV.9 for further detail about the complete inference procedure.
IV.1.2 Configuration model
@@
$P(\bm{k}=\bm{\kappa})=\delta(\bm{\kappa},\bm{\kappa}^{*})\,.$ (39)
@@
IV.1.3 Stochastic block model
The stochastic block model (SBM), in its microcanonical version Peixoto (2017), closely resembles the ER model, where edges are picked uniformly at random. However, unlike the ER model, each node $i$ is associated with a random block $b_{i}\in\{1,2,\dots,B\}$ and instead the number of edges $e_{rs}=\sum_{ij}a_{ij}\delta(b_{i},r)\delta(b_{j},s)$ connecting two blocks $r$ and $s$ is fixed such that there are $E$ edges in total. We summarize the node partition as a tuple $\bm{b}=(b_{i})_{i=1..N}$ and the number of edges between blocks by the edge matrix $\bm{e}=(e_{rs})_{r,s=1..B}$. The blocks are required to be non-empty. Since $B$ is the number of non-empty blocks in $\bm{b}$ and $\bm{e}$ is completely determined by the graph and $\bm{b}$, $\theta=(\bm{b},E)$ are the hyperparameters of the SBM. Consequently, $\bm{b}$ and $E$ must be inferred jointly with the graph. In theory, one could factor the joint prior probability $P(G,E,\bm{b})$ as $P(G|E,\bm{b})P(E)P(\bm{b})$ assuming $E$ and $\bm{b}$ are independent. However, it is more convenient to factor the prior also using $\bm{e}$ and $B$ as intermediate random variables, following